Jenny mixes a 30% saline solution with a 50% saline sloution to make 800 milliliters of a 45% saline solution. How many milliliters of each solution does she use ?
Answer by mananth(16949)

percent ---------------- quantity
Saline I 30 ---------------- x ml
Saline II 50 ------ 800 - x ml
Mixture 45.00% ---------------- 800

30x+50( 800-x) =45*800

30x+40000-50x 36000
30x -50x= 36000 - 40000
-20 x= -4000
/ -20
x=200 ml30.00% Saline I
600ml50.00%Saline II